OpenStudy (hockeygirl10):

Please help I am struggling with this What could you do to increase the electric potential energy between two positively charged particles by a factor of 16?

OpenStudy (anonymous):

What does the electric potential energy depend on? How would you calculate it?

OpenStudy (michele_laino):

the potential energy between two electric charges \(q\) , separated by a distance \(d\) , is: \[U = \frac{{{q^2}}}{d}\] now, if we reduce their separation by a factor \(1/16\), namely, we consder a new separation distance \(d_1=d/16\), then the corresponding potential energy, is: \[{U_1} = \frac{{{q^2}}}{{{d_1}}} = \frac{{{q^2}}}{{d/16}} = ...?\]